OpenStudy (anonymous):

What is the solution of the system? Use substitution. x = –2y + 10 –6y = 6x – 6 z = –3x + 10y A. x = 9, y = –8, z = 114 B. x = –8, y = 9, z = 114 C. x = –8, y = 114, z = 9 D. x = 114, y = 9, z = –8

OpenStudy (dls):

x = –2y + 10 ..(i) –6y = 6x – 6...(ii) z = –3x + 10y... (iii) -6y=6(-2y+10)-6....(iv) Solve for y. Then substitute in (i) to solve for x. Use y and x to solve for z in (iii)

OpenStudy (anonymous):

hope this helps :) http://openstudy.com/study#/updates/4f9309afe4b000ae9ec9d0b4

OpenStudy (anonymous):

it did thx

OpenStudy (anonymous):

No problem :)